# Landau-de Gennes Modelling of Confinement Effects and Cybotactic Clusters in Bent-Core Nematic Liquid Crystals

### v1

@@ IV Confinement Effects for BCN Spatial Equilibria on Square Domains
For example, for the diagonal states D, the paired critical points are D (with co-aligned $\mathbf{n}_{g}$ and $\mathbf{n}_{c}$ in $\Omega_{\parallel\perp}$) and D (with $\mathbf{n}_{g}\cdot\mathbf{n}_{c}=0$ in $\Omega_{\parallel\perp}$) in Fig. 7. The associated $\Omega_{\parallel\perp}$ is located around the square centre in Figure 7. For BD states, the paired critical points are labelled as BD and BD in Fig. 6, using the same reasoning. The corresponding $\Omega_{\parallel\perp}$ is also away from the square edges, the defect lines and localised near the square centre.
Substituting $\sin(2(\theta-\phi))=0$ into the EL equations above, the equations (22)-(23) reduce to
@@
we have $\Delta\theta|_{(0.5,0.5)}=\Delta\phi|_{(0.5,0.5)}=0$ and we deduce that $\Delta\theta|_{\Omega_{\parallel\perp}}\approx 0$ and $\Delta\phi|_{\Omega_{\parallel\perp}}\approx 0$. Similarly, $\theta$ and $\phi$ are almost constant around the square centre for the paired critical points: BD and BD and hence, we deduce that $\Delta\theta|_{\Omega_{\parallel\perp}}\approx 0$ and $\Delta\phi|_{\Omega_{\parallel\perp}}\approx 0$. In other words, the compatibility conditions (24) and (25) are satisfied in $\Omega_{\parallel\perp}$ for the paired critical points in question.
@@ V Numerical methods
The numerical results of Section IV, including the minimisers and critical solutions of the non-dimensionalised free energy in 2D setting $\tilde{F}$ in (14), are calculated by the numerical methods as follows. The study is implemented in Matlab R2021b .
@@ V.1 Spatial discrete method
We use the finite difference method to estimate the spatial derivative on a 2D square $\tilde{\Omega}=[0,1]^{2}$ taking the nodes $(x_{i},y_{j})$, $i,j=0,1,\cdots,N$ with step length $h=\frac{1}{N}$, where
@@ V.2 HiOSD method
We use the HiOSD method [25], which is a generalization of the optimization-based shrinking dimer method [28] to find stable and unstable critical solutions of the Euler-Lagrange equations in (15) on the energy landscape of $\tilde{F}$ in (14).
@@
An index-$k$ critical solution can be found by the $k$-HiOSD dynamics with a certain initial condition. A $k$-HiOSD dynamics is a transformed gradient flow of $\mathbf{x}$ coupling with the search for an orthonormal basis $\mathcal{V}$ which minimises $k$ Rayleigh quotients simultaneously.
@@ V.3 Algorithm for building the connectivity of critical solutions
Following the HiOSD dynamics, we build the connectivity of critical solutions using two algorithms: a downward search that enables us to search for all connected lower-index saddles from an index-$k$ saddle; an upward search with a selected direction to find the higher-index saddles. The two algorithms drive the entire search to navigate up and down on the energy landscape. Please refer to [23] for detailed procedure of downward and upward search.
@@ VI Conclusion
In this paper, we propose and study a generalised LdG-type model for BCN, building on the two-state model proposed by Madhusudhana in [1]. Our model is based on two tensor order parameters: $\mathbf{Q}_{g}$ and $\mathbf{Q}_{c}$ instead of two scalar order parameters, $S_{g}$ and $S_{c}$ as in [1]. Our model has all the capabilities of Madhusudhana’s model and more — it can account for GS and ES directors, spatial inhomogeneities, effects of boundary conditions, geometric frustration, defects etc. Notably, we find that the GS-ES coupling, measured in terms of $D$ in (5), can induce phase transitions at a fixed temperature. For example, we can reproduce the nematic-paranematic phase transition and the paranematic-isotropic phase transitons by reducing $D$, at a fixed high temperature. Madhusudhana reports the same sequence of phase transitions with increasing temperature at fixed $D$, in 3D. We also study phase transitions in 2D although strictly 2D systems are not realistic. However, $\mathbf{Q}_{g}$ and $\mathbf{Q}_{c}$ are reduced to symmetric traceless $2\times 2$ matrices in quasi-2D settings as in Section IV and as such, the 2D phase diagrams in Section III can shed light into planar ordering and disordering as a function of $D$ in such settings. The 2D phase diagrams in Section III certainly help us understand structural phase transitions in confined quasi-2D systems as a function of $D$. Of course, we do not have any physical insight into how to tune or manipulate $D$, except that it is an intrinsic material parameter.
@@
It is commendable that a simple two order parameter model in [1], further generalised in our paper, is capable of capturing some complex experimental trends. This is, in itself, a strong reason for studying such models since they offer tractable approaches to complex multi-physics problem. However, they might miss subtle underpinning physics, important information about molecular shapes and interactions, and can only offer a coarse-grained homogenised perspective. For example, our model can offer no insight into how the opening angle of BCN molecules affects the macroscopic BCN phases. We employ a generic coupling term proportional to $\mathbf{Q}_{g}\cdot\mathbf{Q}_{c}$ as an immediate generalisation of the GS-ES coupling term in [1]. This is a canonical coupling term that arises in homogenisation theory, e.g., when we treat the cybotactic clusters as suspended smectic nanoparticles or nanoclusters in a nematic GS medium, in a dilute limit [16]. However, it is then difficult to justify the elastic energy associated with $\mathbf{Q}_{c}$ or the exact origins of the Ginzburg-Landau type potential for $\mathbf{Q}_{c}$ — what is the underlying physics? Similar coupling terms are also employed in phenomenological studies of ferronematics — dilute suspensions of magnetic nanoparticles in nematic media [15]. Finally, how do we estimate the parameters $M,N$ and $D$ in the free energy (5)? How do we relate these phenomenological parameters to physical experiments on BCNs in a reliable way? Optical measurements probably only yield a mean scalar order parameter as suggested in [1] but one needs quantitative information about $S_{c}$ to infer meaningful information about $M$, NN and $D$. Open questions remain about the interpretability of this model but it is undeniable that this simple two order parameter model captures complex nonlinear phenomena qualitatively and certainly gives insight into cybotactic cluster formation, location and stability in prototype confined systems, making this a worthwhile and interesting study.

### v2

@@ IV Confinement Effects for BCN Spatial Equilibria on Square Domains
In the following, we take the diagonal states D (with co-aligned $\mathbf{n}_{g}$ and $\mathbf{n}_{c}$ in $\Omega_{\parallel\perp}$) and the paired critical state D (with $\mathbf{n}_{g}\cdot\mathbf{n}_{c}=0$ in $\Omega_{\parallel\perp}$) in Fig. 8, the paired critical BD and BD in Fig. 6 as examples.
We first note that, for both types of critical points—those with parallel and those with orthogonal $\mathbf{n}_{g}$ and $\mathbf{n}_{c}$—the conditions (23) and (24) can be satisfied at least on $\Omega_{\parallel\perp}$. For D and D, the associated $\Omega_{\parallel\perp}$ is located around the square centre in Figure 8. For BD and BD, the corresponding $\Omega_{\parallel\perp}$ is also away from the square edges, the defect lines and localised near the square centre. Substituting $\sin(2(\theta-\phi))=0$ into the EL equations above, the equations (23)-(24) reduce to
@@
we have $\Delta\theta|_{(0.5,0.5)}=\Delta\phi|_{(0.5,0.5)}=0$ and we deduce that $\Delta\theta|_{\Omega_{\parallel\perp}}\approx 0$ and $\Delta\phi|_{\Omega_{\parallel\perp}}\approx 0$. Similarly, $\theta$ and $\phi$ are almost constant around the square centre for the paired critical points: BD and BD and hence, we deduce that $\Delta\theta|_{\Omega_{\parallel\perp}}\approx 0$ and $\Delta\phi|_{\Omega_{\parallel\perp}}\approx 0$.
@@
The domains with $\tilde{\lambda}^{2}=2500$ and $5000$ correspond to experimentally relevant micron-scale areas of $(0.26\mu m)^{2}$ and $(0.52\mu m)^{2}$, respectively. As the domain size $\lambda^{2}$ increases, the defects become smaller, requiring a finer computational grid. This, in turn, leads to a higher computational cost, lower numerical accuracy, and greater difficulty in achieving convergence. As discussed in the context of the LdG theory for nematics in [30] and [25], computations on small domains remain relevant for large domains. The solutions of the Euler-Lagrange equations on small domains serve as excellent initial conditions for numerical continuation schemes, for large values of $\lambda^{2}$. Most of the solution branches, for small domains, continue to exist on larger domains although the stability properties might change, i.e., some solution branches lose stability as domain size increases. However, some solution branches retain stability or gain stability, as domain size increases or $\lambda^{2}$ increases. The unstable solutions also play a crucial role in the system dynamics and the selection of stable solutions for multistable systems, and hence, our results for $\lambda^{2}=2500$ (although not large enough for experiments on the several-hundred-micrometer scale), can still provide direct/indirect guidance for future experimental studies.
@@ V Numerical methods
The numerical results of Section IV, including the minimisers and critical points of the non-dimensionalised free energy in 2D setting $\tilde{F}$ in (15), are calculated by the numerical methods as follows. The study is implemented in Matlab R2021b .
@@ V.1 Spatial discrete method
We use the finite difference method to estimate the spatial derivative on a 2D square $\tilde{\Omega}=[0,1]^{2}$ taking the nodes $(x_{i},y_{j})$, $i,j=0,1,\cdots,N_{0}$ with step length $h=\frac{1}{N_{0}}$, where
@@ V.2 HiOSD method
We use the HiOSD method [31], which is a generalization of the optimization-based shrinking dimer method [32] to find stable and unstable solutions of the Euler-Lagrange equations in (16) on the energy landscape of $\tilde{F}$ in (15).
@@
An index-$k$ critical point can be found by the $k$-HiOSD dynamics with a certain initial condition. A $k$-HiOSD dynamics is a transformed gradient flow of $\mathbf{x}$ coupling with the search for an orthonormal basis $\mathcal{V}$ which minimises $k$ Rayleigh quotients simultaneously.
@@ V.3 Algorithm for building the connectivity of critical solutions
Following the HiOSD dynamics, we build the connectivity of critical points using two algorithms: a downward search that enables us to search for all connected lower-index saddles from an index-$k$ saddle; an upward search with a selected direction to find the higher-index saddles. The two algorithms drive the entire search to navigate up and down on the energy landscape. Please refer to [9] for detailed procedure of downward and upward search.
@@ VI Conclusion
In this paper, we propose and study a generalised LdG-type model for BCN, building on the two-state model proposed by Madhusudana in [17]. Our model is based on two tensor order parameters: $\mathbf{Q}_{g}$ and $\mathbf{Q}_{c}$ instead of two scalar order parameters, $S_{g}$ and $S_{c}$ as in [17]. Our model has all the capabilities of Madhusudana’s model and more — it can account for the GS director and additional directors induced by the cybotactic clusters, spatial inhomogeneities, effects of boundary conditions, geometric frustration, defects etc. Notably, we find that the GS-ES coupling, measured in terms of $D$ in (5), can induce phase transitions at a fixed temperature. For example, we can reproduce the nematic-paranematic phase transition and the paranematic-isotropic phase transitions by reducing $D$, at a fixed high temperature. Madhusudana reports the same sequence of phase transitions with increasing temperature at fixed $D$, in 3D. We also study phase transitions in 2D although strictly 2D systems are not realistic. However, $\mathbf{Q}_{g}$ and $\mathbf{Q}_{c}$ are reduced to symmetric traceless $2\times 2$ matrices in quasi-2D settings as in Section IV and as such, the 2D phase diagrams in Section III can shed light into planar ordering and disordering as a function of $D$ in such settings. The 2D phase diagrams in Section III certainly help us understand structural phase transitions in confined quasi-2D systems as a function of $D$. Of course, we do not have any physical insight into how to tune or manipulate $D$, except that it is an intrinsic material parameter.
@@
It is commendable that a simple two order parameter model in [17], further generalised in our paper, is capable of capturing some complex experimental trends. This is, in itself, a strong reason for studying such models since they offer tractable approaches to complex multi-physics problem. However, they might miss subtle underpinning physics about the cybotactic clusters and can only offer a coarse-grained homogenised perspective. For example, our model does not have additional positional order parameters to account for density modulations within cybotactic clusters. This model does not account for the spacing between the cybotactic clusters or the Smectic-type ordering within the cybotactic clusters.To describe these additional features, one may need to adopt a model for SmC‑type ordering, such as [29], in which a density order parameter is introduced in addition to $\mathbf{Q}_{c}$, and a coupling term between the layer normal (which can be represented through the first‑ or second‑order derivatives of the density field) and $\mathbf{Q}_{c}$ is included in the free energy. This leads to a more complex mathematical formulation, which will be explored in our future work. Further, Madhusudana’s model and our generalised model can offer no insight into how the opening angle of BCN molecules affects the macroscopic BCN phases. We employ a generic coupling term proportional to $\mathbf{Q}_{g}\cdot\mathbf{Q}_{c}$ as an immediate generalisation of the GS-ES coupling term in [17]. This is a canonical coupling term that arises in the dilute limit of suspensions, e.g., when we treat the cybotactic clusters as a dilute suspension of smectic nanoparticles or nanoclusters in a nematic GS medium, consistent with the assumptions in Madhusdana’s paper [1]. Similar coupling terms are also employed in phenomenological studies of ferronematics — dilute suspensions of magnetic nanoparticles in nematic media [8]. This coupling term can miss nonlinear phenomena. For example, as the global nematic order $\mathbf{Q}_{g}$ increases, $\mathbf{Q}_{c}$ may exhibit nonlinear behavior such as saturation characteristics. In bent‑core nematics, the molecules are intrinsically bent and typically respond differently along the bending plane and the perpendicular plane. Possible improvements include introducing higher‑order or anisotropic couplings. Further, it is often difficult to decide on the exact form of the elastic energy associated with $\mathbf{Q}_{c}$ or the exact origins of the Ginzburg-Landau type potential for $\mathbf{Q}_{c}$ — what is the underlying physics? Finally, how do we estimate the parameters $M,N$ and $D$ in the free energy and relate these phenomenological parameters to physical experiments on BCNs in a reliable way? Optical measurements probably only yield a mean scalar order parameter as suggested in [17] but one needs quantitative information about $S_{c}$ to infer meaningful information about $M$, NN and $D$. Open questions remain about the interpretability of this model but it is undeniable that this simple two order parameter model captures complex nonlinear phenomena qualitatively and certainly gives insight into cybotactic cluster formation, location and stability in prototype confined systems, making this a worthwhile and interesting study.